# Sample-efficient Imitative Multi-token Decision Transformer for Generalizable Real World Driving

### v1

@@ Abstract
Reinforcement learning via sequence modeling has shown remarkable promise in autonomous systems, harnessing the power of offline datasets to make informed decisions in simulated environments. However, the full potential of such methods in complex dynamic environments remain to be discovered. In autonomous driving domain, learning-based agents face significant challenges when transferring knowledge from simulated to real-world settings and the performance is also significantly impacted by data distribution shift. To address these issue, we propose Sample-efficient Imitative Multi-token Decision Transformer (SimDT). SimDT introduces multi-token prediction, imitative online learning and prioritized experience replay to Decision Transformer. The performance is evaluated through empirical experiments and results exceed popular imitation and reinforcement learning algorithms on Waymax benchmark.
Keywords: Reinforcement Learning, Motion Planning, Autonomous Driving
@@ 1 Introduction
The realm of autonomous driving research has witnessed remarkable progress, with simulation technologies [1][2][3][4] reaching unprecedented levels of realism and the burgeoning availability of real-world driving datasets [5][6][7][8]. Despite these advancements, data-driven planning continues to confront a formidable obstacle: the infinite state space and extensive data distribution characteristic of real-world driving.
@@
Imitation learning approaches encounter hurdles [9][10] when presented with scenarios that deviate from the training distribution, exemplified by rare events like emergency braking for unforeseen obstacles. Similarly, these methods grapple with long-tail distribution phenomena, such as navigating through unexpected weather conditions or handling the erratic movements of a jaywalking pedestrian. On the other hand, reinforcement learning (RL) strategies aim to cultivate policies through reward-based learning. RL has difficulty bridging the sim-real gap and sampling efficiency [11]. It often struggle to extrapolate a driving policy that encapsulates the nuanced decision-making process of an experienced human driver, especially when the simulator lacks interactivity or the scenario falls short of realism [12].
@@
Traditional reinforcement learning approaches also struggle with large state space, long-horizon planning and sparse rewards, which are also characteristic of real-world driving scenarios. Decision Transformer [13] leverages a transformer-based architecture to learn policies for decision-making in reinforcement learning tasks via sequence modeling. Despite its potential on scaling with large state space [14], the original architecture and pipeline is deigned for offline learning and is not enough for complex and dynamic autonomous driving task. Classic RL techniques such as prioritized experience replay [15] which is dealing with large scale dataset cannot naturally be applied as Decision Transformer does not compute temporal-difference.
On the other hand, concentrating solely on single-token prediction renders the model excessively susceptible to immediate contextual patterns, thereby neglecting the necessity for more extensive deliberation over protracted sequences. Models trained through next-token prediction methodologies necessitate substantial dataset to achieve a degree of intelligence that humans attain with considerably less token exposure [16]. Receding Horizon Control[17] is a control method that optimizes decision-making over a rolling time horizon, constantly updating its strategy based on newly acquired information. This approach is analogous to multi-token prediction in decision transformers and has potential shifting from myopic to panoramic prediction closed to human cognitive processes.
This paper seeks to address these challenges by proposing an improved Decision Transformer network and a hybrid learning framework that leverages the complementary strengths of imitation and reinforcement learning. Experiment results indicate that our approach yields a substantial enhancement in performance with improvements observed in terms of policy robustness and sample efficiency. The main contributions are as follows:
@@
  • We present a fully online imitative Decision Transformer pipeline designed for wide data distribution across large-scale real-world driving dataset.
@@
  • We propose multi-token Decision Transformer architecture for receding horizon control to enhance long-horizon prediction and broaden attention field.
@@
  • We introduce prioritized experience replay to Decision Transformer and enables sample-efficient training for large-scale sequence modelling based reinforcement learning.
@@ 2 Related Work
Reinforcement learning via sequence modeling. Trajectory Transformer [18] and Decision Transformer (DT) [13] are pioneer in this area, leveraging transformer architectures to model sequences of state-action-reward trajectories and predicting future actions in offline manner. Following work [19] [20] [21] [22] extends leverage the power of transformers for efficient and generalized decision-making in RL. Online DT [23] and Hyper DT [24] adapt original concept for online settings and interacts with environments. However, previous work are done on relatively simple environments compared to autonomous driving environment.
Multi-token prediction. Transformers have significantly impacted NLP since their inception [25], outperforming RNNs and LSTMs by processing sequences in parallel and efficiently handling long-range dependencies. Subsequent models like GPT [26] and BERT [27] have refined the architecture, enhancing pre-training, fine-tuning, and scalability. Recent studies explore multi-token prediction on semantic representation [28], streamline computation [29], prediction technique [30] and multilingual [31]. However,Focusing only on single-token prediction makes the model too sensitive to immediate context and overlooks the need for deeper analysis of longer sequences[16]. This paper extends the concept to Decision Transformer and explore the potential benefits of multi-token prediction for motion planning.
Learning with real-world driving data. Many work has been done to accommodate with real-world driving data [5][2][6][7] for generalizable driving policy. Lu et al. [32] explores the cooperation between reinforcement learning and imitation learning for real-world driving data. DriveIRL [33] designs an inverse reinforcement learning architecture to learn score component in complex heavy traffic scenarios. TuPlan [9] combines both learning methods with rule-based method for real-world planning. Trajeglish [10] models real world traffic auto-repressively as language processing problem and it has most similar network structure as MSDT. Our approach differs as we apply imitative reinforcement learning for pretraining and online real-world planning adaption for performance enhancement.
@@ 3 Methods
In this section, we introduce MSDT, a multi-token sample-efficient reinforcement learning framework via sequence modelling for dynamic driving scenarios. MSDT consists of three components: multi-token decision transformer, online imitative pipeline and prioritized experience replay.
Figure 2: The general network architecture of SimDT. Feature encoding is applied to extract complex real-world driving perceptual data to small but meaningful embeddings. Inside causal transformer, attention relationship are calculated for past context length of $[(s_{t},a_{t},g_{t})]_{T-c}^{T}$. The decoder now predicts multi-tokens for actions and only the first action is applied during the inference stage.
@@ 3.1 Network Structure
Since the real-world driving environment is complex and dynamic, specific feature encoding network is designed for the states representation. Real-world driving state contains many perceptual information such as obstacles, road map, traffic and so on. We follow the vectorized representation to organize road map as polylines and then extract with Polyline Encoder [34]. Obstacle with past 10 historical information are recorded in terms of $\bm{[}p_{x},p_{y},v_{x},v_{y},l,w]$. Obstacle and traffic embedding are extracted with multi-layer perception network.
The work further extends the method to goal-conditioned reinforcement learning by adding the relative vector distance between ego vehicle and destination. The importance of goal-condition lies in its influence on the decision-making process of the autonomous agent. Even in an identical environment, the actions taken by the vehicle can vary significantly depending on the specified goal.
Multi-token prediction in causal transformer simultaneously generates multiple tokens in a single forward pass, while still respecting the autoregressive property that ensures each prediction only depends on previously generated tokens. This is typically achieved by using a masked self-attention mechanism that allows the model to consider multiple future positions without violating causal dependencies. As shown in Equation (1), next-token prediction has loss function $L_{a}$ defined as the negative log-likelihood of the policy.
@@
The loss function is modified for multi-token prediction and assume network predict next 3 tokens. Where $\alpha$ and $\beta$ are the coefficient designed for network to learn more about current step action predictions.
$L_{ma}=$ $-\log\pi_{\theta}(a_{t}\mid s_{t:t-c},a_{t-1:t-c},g_{t:t-c})$ (2)
$-\alpha*\log\pi_{\theta}(a_{t+1}\mid s_{t:t-c},a_{t-1:t-c},g_{t:t-c})-\beta*\log\pi_{\theta}(a_{t+2}\mid s_{t:t-c},a_{t-1:t-c},g_{t:t-c})$
3.2 Online Imitative Training Pipeline
Algorithm 1 Online Imitative Training Pipeline
@@ 3.2 Online Imitative Training Pipeline
     HindsightReturnRelabeling: $D_{traj}\leftarrow D_{trans}$, $[[(s_{i,j},a_{i,j},g_{i,j})]_{i=0}^{T}]_{j=0}^{A/T}\leftarrow\left[(s_{i},a_{i},r_{i})\right]_{i=0}^{A}$
@@
The general idea of the proposed algorithm is to perform sample-efficient online imitative reinforcement learning with off-policy expert data for pre-training at beginning. Subsequently, the model undergoes a mixed on-policy adaptation phase which is introduced at the mid-point of the training process. The core concept behind is to quickly shift the distribution towards the expert behavior at beginning and reduce environmental distribution shift with on-policy adaption. Note online adaption and imitative reinforcement learning are performed concurrently after mid of training, this helps the network no to fall into online local minimal.
Imitative reinforcement learning is done by applying similar concept as Shaped IL [35] and GRI [36] where reward is shaped for expert demonstration data. Following same implementation in [32], expert data from real world driving trajectory was converted to expert agent actions with inverse kinematics. We also design negative reward for offroad and overlap (collision) behavior. The network will learn good behavior through imitation reward and bad actions through online interaction with offraod and overlap rewards. The overall online imitative pipeline is essential to achieve the greater data-distributed policy described in Figure 1.
@@
$R_{imitaiton}=\begin{cases}1.0&\text{if }\text{{log\_divergence}}<0.2,\\ 0.0&\text{if }\text{{log\_divergence}}>0.2.\end{cases}$ (3)
$R_{offroad}=-2$ (4)
@@
However, the real-time collected transition level replay buffer does not contain return-to-go as it can only be calculated after episode is finished and all rewards is collected. Similar to Online Decision Transformer, the transition level replay buffer converted to hindsight trajectory replay buffer when fixed amount of trajectories are collected.
@@ 3.3 Prioritized Experience Replay for Decision Transformer
     Execute lines from Algorithm 1 $\triangleright$ Online Data Collection
     HindsightReturnRelabeling: $D_{traj}\leftarrow D_{trans}$, $[[(s_{i,j},a_{i,j},g_{i,j})]_{i=0}^{T}]_{j=0}^{A/T}\leftarrow\left[(s_{i},a_{i},r_{i})\right]_{i=0}^{A}$
@@
         $D_{single}^{per}\leftarrow\{[[(s_{i,j},a_{i,j},g_{i,j})]_{t-c}^{t}]_{j},L_{single}\}$
         $D_{overall}^{per}\leftarrow\{[[(s_{i,j},a_{i,j},g_{i,j})]_{t-c}^{t}]_{j},L_{overall}\}$
@@
Prioritized Experience Replay (PER) selectively samples experiences with high temporal-difference errors from the replay buffer for focusing on more informative experiences. However, the Decision Transformer doesn’t use temporal-difference errors, precluding direct application of PER. Instead, we adapt by using action loss to gauge transition importance within the Decision Transformer, which assesses state-action-return relationships. The design concept is that if the model’s predicted actions diverge from actual ones, it indicates a misinterpretation of the environment.
On top of above architecture, extra replay buffers are designed to store prioritised sampled trajectories based on action loss. The action loss represents the difference between the actions predicted by the policy network and the actual actions taken. A low actor loss means that the policy network’s predictions are close to the actual actions, while a high actor loss means that the predictions are far from the actual actions. Prioritised sampled trajectories are stroed based on following criteria:
Criterion 1: Preservation of transitions with maximal single-step action discrepancy: This methodology concentrates on isolating the instances wherein the model’s prognostications manifest the greatest deviation from expected accuracy. Such a strategy is instrumental in directing the model’s learning efforts towards ameliorating its most significant errors.
Criterion 2: Preservation of transitions with maximal cumulative action discrepancy: This methodology is characterized by its emphasis on identifying and retaining sequences wherein the aggregate error of the model’s predictions reaches its apex. This approach holds particular utility for endeavors aimed at refining the model’s performance across a continuum of actions.
The replay buffers store data based on high value in low value out. The prioritized experience replay buffer is sampled for training every fixed amount of episode and its priorities are updated at meantime. The goal for the proposed prioritized experience replay in this paper is to prioritize the trajectories where model has biggest misunderstanding of the corresponding scenarios, and therefore to prioritize on long-tail scenarios.

### v2

# Sample-efficient Imitative Multi-token Decision Transformer for Real-world Driving

@@ Abstract
Recent advancements in autonomous driving technologies involve the capability to effectively process and learn from extensive real-world driving data. Current imitation learning and offline reinforcement learning methods have shown remarkable promise in autonomous systems, harnessing the power of offline datasets to make informed decisions in open-loop (non-reactive agents) settings. However, learning-based agents face significant challenges when transferring knowledge from open-loop to closed-loop (reactive agents) environment. The performance is significantly impacted by data distribution shift, sample efficiency, the complexity of uncovering hidden world models and physics. To address these issues, we propose Sample-efficient Imitative Multi-token Decision Transformer (SimDT). SimDT introduces multi-token prediction, online imitative learning pipeline and prioritized experience replay to sequence-modelling reinforcement learning. The performance is evaluated through empirical experiments and results exceed popular imitation and reinforcement learning algorithms both in open-loop and closed-loop settings on Waymax benchmark. SimDT exhibits 41% reduction in collision rate and 18% improvement in reaching the destination compared with the baseline method.
@@ 1 Introduction
The rapid advancement in machine learning has led to unprecedented achievements in the autonomous systems domain. One of the critical methodologies in training these neural networks involves leveraging large offline datasets. These datasets provide the foundational knowledge that neural networks require to learn and make predictions or decisions. However, deploying these offline-trained neural networks in real-world scenarios or simulators with reactive agents presents challenges stemming from the distribution shift between the dataset and the application environment.
@@
Learning from Demonstration (Dauner et al., 2023) (Cheng et al., 2023) (Huang et al., 2023) (Levine et al., 2020) encounter hurdles when presented scenarios deviate from the training distribution, exemplified by rare events like emergency braking for unforeseen obstacles. Similarly, these methods grapple with long-tail distribution phenomena during closed-loop tests, such as navigating through unexpected weather conditions or handling the erratic movements of jaywalking pedestrians.
@@
Interacting with environments allows for improved management of out-of-distribution issues in closed-loop tests by learning policies focused on reward maximization, thus facilitating adaptation. However, reinforcement learning (RL) faces challenges in overcoming the simulation-reality discrepancy and suffers from low sampling efficiency (Kiran et al., 2022) (Liu et al., 2022). Traditional reinforcement learning methods also grapple with the complexities of large state spaces, long-term planning, and sparse rewards, mirroring the challenges of real-world driving scenarios. Decision Transformer (Chen et al., 2021b) leverages a transformer-based architecture to learn policies for decision-making in reinforcement learning tasks via sequence modeling. Despite its potential to scale with large state space (Zhou et al., 2024), the original architecture and pipeline are designed for offline learning and are not enough for complex and dynamic autonomous driving tasks. Classic RL techniques like prioritized experience replay (Schaul et al., 2016) are designed for large-scale datasets but cannot be applied as the Decision Transformer does not compute temporal differences.
On the other hand, single-token prediction does not fully fit the nature of autonomous driving and has difficulty uncovering the hidden physical world model. Concentrating solely on single-token prediction renders the model excessively susceptible to immediate contextual patterns, thereby neglecting the necessity for more extensive deliberation over protracted sequences. Models developed through next-token prediction techniques necessitate extensive datasets to attain a level of cognitive proficiency that humans achieve with significantly less exposure to tokens (Gloeckle et al., 2024).
This paper seeks to address the challenges associated with closed-loop autonomous driving by introducing Sample-efficient Imitative Multi-token Decision Transformer. The proposed approach leverages real-world driving data and realistic simulators, providing scenarios that are both realistic and conducive to online adaptation to distribution shifts. By employing multi-token prediction via sequence modeling of (state, action, return_to_go) pairs, the network is enabled to discern the quality of various action sequences, thereby gaining a deeper insight into the underlying world model. Experiment results indicate that our approach yields a substantial performance enhancement in terms of policy robustness and sample efficiency both in open-loop and closed-loop settings. SimDT exhibits 41% reduction in collision rate and 18% improvement in reaching the destination compared with the baseline method. The work is conducted entirely with Jax, which facilitates highly efficient training on large-scale data and real-time inference. The inference time is 1.63 milliseconds for SimDT(median) on RTX 3090. The main contributions are as follows:
@@
  • We present an online imitative reinforcement learning pipeline designed for wide data distribution across a collaboration of real-world driving datasets and simulators.
@@
  • We propose multi-token Decision Transformer architecture for receding horizon control to enhance long-horizon prediction and broaden the attention field.
@@
  • We introduce prioritized experience replay to sequence modeling-based reinforcement learning and enable sample-efficient training for large-scale data.
@@ 2 Related Work
Reinforcement learning via sequence modeling. Trajectory Transformer (Janner et al., 2021) and Decision Transformer (DT) (Chen et al., 2021b) are pioneers in this area, leveraging transformer architectures to model sequences of state-action-reward trajectories and predicting future actions in an offline manner. Following work (Lee et al., 2022) Meng et al. (2022) (Wu et al., 2023) (Badrinath et al., 2023) extends leverage the power of transformers for efficient and generalized decision-making in RL. Online DT (Zheng et al., 2022) and Hyper DT (Xu et al., 2023) adapt the original concept for online settings and interact with environments. However, previous work are done on relatively simple environments compared to autonomous driving environments.
Multi-token prediction. Transformers have significantly impacted NLP since their inception (Vaswani et al., 2017), outperforming RNNs and LSTMs by processing sequences in parallel and efficiently handling long-range dependencies. Subsequent models like GPT (Radford et al., 2018) and BERT (Devlin et al., 2019) have refined the architecture, enhancing pre-training, fine-tuning, and scalability. Recent studies explore multi-token prediction on semantic representation (Kitaev et al., 2020), streamline computation (Wang & Cho, 2019), prediction technique (Qi et al., 2020) and multilingual (Jiang et al., 2020). However, focusing only on single-token prediction makes the model too sensitive to the immediate context and overlooks the need for deeper analysis of longer sequences(Gloeckle et al., 2024). This paper extends the concept to Decision Transformer and explores the potential benefits of multi-token prediction for motion planning.
Learning with real-world driving data. Much work has been done to accommodate with real-world driving data (Sun et al., 2020)(H. Caesar, 2021)(Houston et al., 2020)(Phan-Minh et al., 2023) for generalizable driving policy. Lu et al. (2023) explores the cooperation between reinforcement learning and imitation learning in terms of loss design for real-world driving data. TuPlan (Dauner et al., 2023) combines both learning methods with rule-based methods for real-world planning. Guided Online Distillation (Li et al., 2023) adapts real-world data for reinforcement learning online distillation in MetaDrive Simulator (Li et al., 2022). Trajeglish (Philion et al., 2024) models real-world traffic autoregressively as a language processing problem with a causal transformer. Our approach is most similar to Trajeglish as both works tokenize expert driving logs to state-action sequences, encode agent and map information for better scene understanding and finally output actions for agent control. However, Trajeglish is fundamentally a supervised learning method and it inevitably has training and test distribution mismatch. Our approach differs as SimDT additionally models return to state-action sequences and convert supervised learning to reinforcement learning sequence modeling (state-action-return) problem to further solve the distribution shift issue. What’s more, SimDT further develops online adaptation and multi-token prediction in contrast to Trajeglish’s next-token prediction to further enhance its modeling of the driving data.
@@ 3 Methods
In this section, we introduce SimDT, a novel sample-efficient reinforcement learning framework designed for dynamic driving scenarios through sequence modeling. SimDT integrates a multi-token decision transformer for nuanced decision-making, an online imitative reinforcement learning pipeline for continuous adaptation and improvement, and a prioritized experience replay mechanism to enhance learning efficiency by focusing on more informative experiences.
@@ 3.1 Network Structure
Since the real-world driving environment is complex and dynamic, a specific feature encoding network is designed for the state representation. The real-world driving state contains perceptual information such as obstacles, road map, traffic and so on. We follow the vectorized representation to organize the road map as polylines and then extract with Polyline Encoder (Shi et al., 2022). Obstacles with past 10 historical information are recorded in terms of [position_x, position_y, yaw, speed, length, width]. Obstacle and traffic embedding are extracted with multi-layer perception network.
The work further extends the method to goal-conditioned reinforcement learning by adding the relative vector goal between the ego vehicle and the destination in terms of ($x,y$). The importance of goal condition lies in its influence on the decision-making process of the autonomous agent. Even in an identical environment, the actions taken by the vehicle can vary significantly depending on the specified goal position.
Real-world data are recorded in terms of trajectory and it is converted to actions via inverse kinematics. Multi-token prediction is designed to enhance the network’s understanding of reference trajectories and short-term momentum impact by learning a fixed horizon of reference actions (Figure 2,7). Multi-token prediction in a causal transformer simultaneously generates multiple actions in a single forward pass, while still respecting the autoregressive property that ensures each prediction only depends on previously generated tokens. This is typically achieved by using a masked self-attention mechanism that allows the model to consider multiple future positions without violating causal dependencies. As shown in  equation 1, the next-token prediction has loss function $L_{a}$ defined as the negative log-likelihood of the policy.
@@
The loss function is modified for the next n token prediction. Where $\alpha$, $\beta$ and $\gamma$ are the coefficients designed for the network to learn more about current step action predictions.
$L_{ma}=$ $-\alpha*\log\pi_{\theta}(a_{t}\mid s_{t:t-c},a_{t-1:t-c},g_{t:t-c})$ (2)
$-\beta*\log\pi_{\theta}(a_{t+1}\mid s_{t:t-c},a_{t-1:t-c},g_{t:t-c})-\gamma*\log\pi_{\theta}(a_{t+2}\mid s_{t:t-c},a_{t-1:t-c},g_{t:t-c})$
$-\quad\cdots\cdots\quad-\omega*\log\pi_{\theta}(a_{t+n}\mid s_{t:t-c},a_{t-1:t-c},g_{t:t-c})$
Figure 2: The general network architecture of SimDT. Feature encoding is applied to extract complex real-world driving perceptual data to small but meaningful embeddings. Inside the causal transformer, attention relationships are calculated for past context length of $[(s_{t},a_{t},g_{t})]_{T-c}^{T}$. The decoder now predicts multi-tokens for actions and only the first action is applied during the inference stage.
3.2 Online Imitative Reinforcement Learning Pipeline
The general idea of the proposed Algorithm 1 is to perform online imitative reinforcement learning with off-policy expert data for pre-training at the beginning. Subsequently, the model undergoes a mixed on-policy adaptation phase which is introduced at the mid-point of the training process. The core concept behind this is to quickly shift the distribution towards the expert behavior at the beginning and reduce environmental distribution shift with on-policy adaptation (OPA). Note online policy adaptation and imitative reinforcement learning are performed concurrently after mid of training, this helps the network not to fall into online local minimal.
Imitative reinforcement learning is done by applying a similar concept as Shaped IL (Judah et al., 2014) and GRI (Chen et al., 2021a) where the reward is shaped for expert demonstration data. Following the same implementation in (Lu et al., 2023), expert data from real-world driving trajectory was converted to expert agent actions with inverse kinematics. We also design negative rewards for off-road and overlap (collision) behavior. The network learns good behavior through imitation rewards and bad actions through online interaction with off-road and overlap rewards. The overall online imitative reinforcement pipeline is essential to achieve the greater data-distributed policy described in Figure 1.
Algorithm 1 Online Imitative Reinforcement Learning Pipeline
@@ 3.2 Online Imitative Reinforcement Learning Pipeline
     HindsightReturnRelabeling: $D_{traj}\leftarrow D_{trans}$equiv $[[(s_{i,j},a_{i,j},g_{i,j})]_{i=0}^{T}]_{j=0}^{A/T}\leftarrow\left[(s_{i},a_{i},r_{i})\right]_{i=0}^{A}$
@@
$R_{imitaiton}=\begin{cases}1.0&\text{if }\text{{log\_divergence}}<0.2,\\ 0.0&\text{if }\text{{log\_divergence}}>0.2,\end{cases}$ (3)
$R_{off-road}=-2$ (4)
@@
where log_divergence is the euclidean distance between the log history of the ego vehicle and that of the controlled agent.
The real-time collected transition level replay buffer does not contain return-to-go as it can only be calculated after the episode is finished and all rewards are collected. Similar to Online DT(Zheng et al., 2022), the transition level replay buffer is converted to hindsight trajectory replay buffer when a fixed amount of trajectories are collected (HindsightReturnRelabeling). Moreover, the dataloader randomly shuffles obstacles in order within the state (ShuffleObstacleOrder) to enhance data augmentation.
@@ 3.3 Prioritized Experience Replay for Decision Transformer
     Execute online data collection from Algorithm 1 $\triangleright$ Online Data Collection
     HindsightReturnRelabeling: $D_{traj}\leftarrow D_{trans}$equiv $[[(s_{i,j},a_{i,j},g_{i,j})]_{i=0}^{T}]_{j=0}^{A/T}\leftarrow\left[(s_{i},a_{i},r_{i})\right]_{i=0}^{A}$
@@
         $D_{single}^{per}\leftarrow\{[[(s_{i,j},a_{i,j},g_{i,j})]_{i=t-c}^{t}]_{j},L_{single}\}$
         $D_{overall}^{per}\leftarrow\{[[(s_{i,j},a_{i,j},g_{i,j})]_{i=t-c}^{t}]_{j},L_{overall}\}$
@@
The proposed imitative reinforcement learning method not only handles large volumes of offline demonstration data but also obtains an infinite amount of data through online interactions. The ability to utilize prioritized experience replay (PER) becomes critical for enhancing sample efficiency. The original proposed PER(Schaul et al., 2016) selectively samples experiences with high temporal-difference errors from the replay buffer for focusing on more informative experiences. However, DT does not use temporal-difference errors and therefore precludes direct application of PER. Instead, we adapt by using action loss to gauge transition importance within the Decision Transformer, which assesses state-action-return relationships. The design concept is that if the model’s predicted actions diverge from actual ones, it indicates a misinterpretation of the environment.
We develop prioritized experience replay for Decision Transformer on top of the previous online imitative training pipeline(Algorithm 2). Extra replay buffers are designed to store prioritized sampled trajectories based on action loss. The action loss represents the difference between the actions predicted by the policy network and the actual actions taken. A low actor loss means that the policy network’s predictions are close to the actual actions, while a high actor loss means that the predictions are far from the actual actions. Different from bootstrapping RL algorithms which sample one pair of state, action and reward $(s_{i},a_{i},r_{i})$, Decision Transformer samples a context length($c$) of state-action-return pairs $[[(s_{i,j},a_{i,j},g_{i,j})]_{i=t-c}^{t}]_{j}$ for sequence modeling. Therefore, the replay buffers designed for the decision transformer store a set of state-action-return pairs with a specified context length. Our proposed prioritization rule takes advantage of these state-action-return pairs, which try to analyze and understand either single scene or cumulative scenes. The proposed method prioritizes trajectories based on the following criteria:
Criterion 1: Preservation of transitions which contains single-step action discrepancy($L_{single}$): This methodology concentrates on isolating the instances where the model’s prognostications manifest the greatest deviation from expected accuracy. Such a strategy is instrumental in directing the model’s learning efforts toward ameliorating its most significant errors. The replay buffer for single-step action discrepancy $D_{single}^{per}$ can generally hold unexpected collisions and out-of-line actions.
Criterion 2: Preservation of transitions with maximal cumulative action discrepancy($L_{overall}$): This methodology is characterized by its emphasis on identifying and retaining sequences wherein the aggregate error of the model’s predictions reaches its apex. It holds particular utility for endeavors aimed at refining the model’s performance across a continuum of actions for context length of state-action-return pairs. The replay buffer for cumulative action discrepancy $D_{overall}^{per}$ is useful for preserving rare or long-tail scenarios.
The two replay buffers store data based on high value in low value out. The prioritized experience replay buffer is sampled for training every fixed amount of episode and its priorities are updated in the meantime. The goal for the proposed prioritized experience replay in this paper is to prioritize the trajectories where the model has the biggest misunderstanding of the corresponding scenarios, and therefore to prioritize long-tail scenarios. The final training pipeline is illustrated in Figure 6.